Hello my name is kevin and im an exhange student from spain can you please explain to me the process of how to solve these by graphing?
{2x+9=y
{y=-4x-3

Answer by Edwin McCravy(20086)

{2x+9=y
{y=-4x-3

Get some point of each line.

Get some arbitrary points on the line for the first equation

2x + 9 = y            |2x + 9 = y             |2x + 9 = y
Let x = 0, then       |Let x = -4, then       |Let x = -8, then   
2(0) + 9 = y          |2(-4) + 9 = y          |2(-8) + 9 = y      
       9 = y          |  -8  + 9 = y          |  -16 + 9 = y    
                      |        1 = y          |      -7 = y  
So one point is (0,9) |So one point is (-4,1) |So one point is (-8,-7)

Let's plot those three points:

 

Let's draw a line through them.

 

--------------------------

{y=-4x-3

Get some arbitrary points on the line for the second equation

y = -4x - 3           |y = -4x - 3            |y = -4x - 3           
Let x = 0, then       |Let x = 1, then        |Let x = -3, then   
y = -4(0) - 3         |y = -4(1) - 3          |y = -4(-3) - 3      
y = 0 - 3             |y = -4 - 3             |y = 12 - 3    
y = -3                |y = -7                 |y = 9  
So one point is (0,-3)|So one point is (1,-7) |So one point is (-3,9)

Let's plot those three points:

 

Let's draw a line through them.

 

Maybe you can tell what that point of intersection is just by
looking. If you can't, draw a vertical line straight down from
that point to the x-axis and one horizontal over to the y-axis:



The vertical line intersects the x axis at -2
The horizontal line intersects the y-axis at 5.

So the point of intersection is (-2,5)

The solution is x = -2 and y = 5

Edwin